OpenStudy (anonymous):

How many different three-ingredient sandwiches can be made if 8 ingredients are available?

OpenStudy (experimentx):

8*7*5

OpenStudy (anonymous):

Thank you (:

OpenStudy (experimentx):

did you understand what did i do??

OpenStudy (anonymous):

Yeah, pretty much. You did the permutations. But does the 6 reduce down to a 2? and then cancel? that's the only thing that got me.

OpenStudy (anonymous):

I meant reduce down to a 3 lol now I think I'm just confusing myself.

OpenStudy (anonymous):

experiment I'd love to know what you just did.

OpenStudy (inkyvoyd):

How many different three-ingredient sandwiches can be made if 8 ingredients are available?

OpenStudy (inkyvoyd):

Now, we are using nCr because order does not matter.

OpenStudy (inkyvoyd):

cheese, tomato, lettuce, is the same as tomato, cheese, lettuce, right?

OpenStudy (inkyvoyd):

nCr=(nPr)/r! nPr=(n!)/(n-r)! nCr=(n!)/[r!(n-r)!]

OpenStudy (inkyvoyd):

n=8 r=3

OpenStudy (inkyvoyd):

8!/[3!(8-3)!] =8!/(3!5!) =(8*7*6)/3! =8*7

OpenStudy (anonymous):

I honestly just don't understand how to use combinations or permutations. I'm a bit better at permuatations than combinations, but both confuse me. But I'm trying to follow through with what you are saying, so thank you lol.

OpenStudy (inkyvoyd):

Alright, let me explain just exactly what combinations and permutations are

OpenStudy (inkyvoyd):

Well, you understand permutations, right?

OpenStudy (anonymous):

yeah, I'm pretty sure I do.

OpenStudy (inkyvoyd):

ok. Now, there is only a small difference between combinations and permutations.

OpenStudy (inkyvoyd):

in permutations, "order" matters.

OpenStudy (inkyvoyd):

Say, the winning number of a lottery ticket.

OpenStudy (anonymous):

right

OpenStudy (inkyvoyd):

1119 is different from 9111

OpenStudy (inkyvoyd):

in combinations, however "order" does not matter.

OpenStudy (inkyvoyd):

if we were to take 2 random people from a group of 10 people, selecting Bob then Joe is the same as selecting Joe then Bob

OpenStudy (inkyvoyd):

Now, there are formulas for permuations and combinations, and unfortunately, these require memorization.

OpenStudy (inkyvoyd):

(they don't if you understand them intuitively, however)

OpenStudy (anonymous):

alright, so I just have to memorize the permutations and combinations formula and try and learn how to do them?

OpenStudy (inkyvoyd):

Well, you can try to understand why they are what they are, and then try to memorize; that helps.

OpenStudy (inkyvoyd):

I'll try to explain...

OpenStudy (anonymous):

Yeah I can do that. That's a good idea, thank you. But you don't have to explain it to me if you don't want to. haha but I'd appreciate if you did. :)

OpenStudy (inkyvoyd):

Ok, let's say we have a 8 digit password (0-9), and the numbers cannot be repeated.

OpenStudy (inkyvoyd):

In other words, that would be 10 choices for the first, 9 for the second (because one is taken up by the first), 8 for the third, 7 for the fourth, 6 for the fifth, 5 for the sixth, 4 for the seventh, and 3 for the eighth.

OpenStudy (inkyvoyd):

Now we multiply those all out.

OpenStudy (inkyvoyd):

10*9*8*7*6*5*4*3 I mean, multiply those by each other.

OpenStudy (anonymous):

yes, so that would be 1,814,400 right?

OpenStudy (inkyvoyd):

Not sure. But, what's important is to notice it's striking resemblance to the factorial function. The two must be related.

OpenStudy (inkyvoyd):

in fact, we didn't select the last two digits for the password, right?

OpenStudy (anonymous):

right

OpenStudy (inkyvoyd):

those last two digits happen to be 2!

OpenStudy (inkyvoyd):

(*2*1)

OpenStudy (inkyvoyd):

So, we literally just divided 10! by 2 factorial, right?

OpenStudy (anonymous):

I think so. Like do you meant we divided them before we found out what to multiply?

OpenStudy (inkyvoyd):

Well, we multiplied down to 3 (starting from 10)

OpenStudy (inkyvoyd):

We didn't multiply by 2 and by 1, because those were the ninth and tenth digits, and we wer eonly looking for the first eight.

OpenStudy (anonymous):

Oh I see, that helps me understand it a bit more. So does this "rule" follow for both permutations and combinations?

OpenStudy (inkyvoyd):

We can represent this mathematically.

OpenStudy (inkyvoyd):

In general, if you have n cases to choose from, and choose r different cases

OpenStudy (inkyvoyd):

nPr is shorthand for writing the permuation

OpenStudy (inkyvoyd):

It's shorthand, but we actually express nPr as n!/(n-r)!

OpenStudy (inkyvoyd):

the n! in my example would be 10, and the (n-r)! would be (10-8)!, or 2!

OpenStudy (inkyvoyd):

Combinations are slightly harder to explain.

OpenStudy (anonymous):

Oh, I see. I'd imagine Combinations would be harder to explain. Because I'm not good at combinations at all haha. But I understand permutations a bit more now.

OpenStudy (inkyvoyd):

That's good. My head is feeling a little funny, and I'm sort of sleepy, so I'll just tell you the formulas,and how I remember them, as well as when to use which.

OpenStudy (anonymous):

Alright, thank you so much.

OpenStudy (inkyvoyd):

nPr=n!/(n-r)! nCr=nPr/r! Permutation: order matters - common examples -> lottery ticket, password, combination lock (yes, order matters on these, despite the confusing name) Combination: order does not matter - common exmaples -> selecting a group of people out of a larger group, putting stuff in a sandwich (lettuce and tomatos is the same as tomatoes and lettuce)

OpenStudy (inkyvoyd):

If you need more help, just ask more questions on Open study.

OpenStudy (anonymous):

This was awesome Ink, seriously.

OpenStudy (anonymous):

Thank you!

OpenStudy (inkyvoyd):

You're welcome :)

OpenStudy (inkyvoyd):

The only difference formula wise of combinations and permuations is that nCr=nPr/r!

OpenStudy (anonymous):

Alright, I understand them a bit better thanks to you :) thank you for taking your time to help me with something I should have already got down. haha

OpenStudy (experimentx):

Sorry for wrong answer, Lost connection right after that. 9*8*7/3!

OpenStudy (anonymous):

it's alright :)